# Think Together and Work Better: Combining Humans’ and LLMs’ Think-Aloud Outcomes for Effective Text Evaluation

### v1

@@ 3.3.2. TA by LLMs
The LLMs include four different types of representative agent: GPT-4 (Achiam et al., 2023), Llama-3.1-8B-Instruct  [^6] https://llama.meta.com/, Claude-3.5-Sonnet  [^7] https://www.anthropic.com/news/claude-3-5-sonnet, and Gemini-1.5-Pro (Reid et al., 2024). The LLMs simulate a human-like TA process using the CTA method, processing text attributes needed for scoring each dimension in real-time. For this, we construct a prompt template with the identical materials given to humans during the TA: four news articles and summaries, corresponding scores, and dimension definitions. During the TA, we assign a persona (assessor) to the LLMs and make them read the materials and verbalize their thoughts on what to keep in mind when evaluating a targeting dimension. In addition, they are required not to be too much oriented to specific samples (See Appendix A for the prompt templates used in the LLMs’ TA).
@@ 3.4. Checklist Construction
In this stage, we merge the attributes collected from the previous TA process (Section 3.3) of both humans and LLMs. Then, we build the checklists using the combined text attributes. The checklist construction process includes four steps: 1) Component Extraction, 2) Attributes Clustering, 3) Question Generation, and 4) Question Validation. Each step can utilize an LLM, such as GPT series, and in our study, we use either GPT-3.5-turbo or GPT-4. The integration of attributes from both sources (humans and LLMs) is expected to create a synergistic effect, where the strengths of human expertise and LLMs complement each other, leading to a more refined and effective checklist (see Appendix B to F for the prompt templates used in each step). A key difference between InteractEval and CheckEval (Lee et al., 2024b) is how components are created. In CheckEval, humans (authors) manually derive components from a rubric. In contrast, InteractEval generates components from a combination of text attributes based on the TA outcomes of both humans (experts) and LLMs, using rubrics and sample texts. This approach reduces bias from relying on a single source and allows for more detailed checklists base on the fine-grained attributes (Pantcheva, 2023; Blattberg and Hoch, 1990; Brindley, 1998; Turner and Upshur, 2002).
@@ 3.4.1. Component extraction
The component extraction process identifies key themes related to a rating dimension from the text attributes generated in the prior TA process. For instance, a component related to coherence may involve attributes such as ”Structure”, ”Logical Flow” and ”Tone style”. This step narrows down the focus areas, laying the groundwork for generating key questions tied to these components. Using an LLM as an extractor, principle components are extracted by prioritizing the most pertinent attributes from both human experts and the LLMs. To avoid noise and redundancy, we prompt the extractor to eliminate ambiguous or repetitive components from the combined attributes. Also, we let the agent combine related attributes into cohesive entities. The number of components is limited to five or fewer to prevent excessive subdivision, which can lead to overlapping meanings and create noise. These components then guide the subsequent steps in checklist development.
@@ 3.4.2. Attributes clustering
After extracting the components from the combined group of attributes, an LLM clusters each attribute under the appropriate component to identify specific associated attributes. This clustering divides the dimension into segmented sections, leveraging the ideas of both humans and LLMs. Each section focuses on a specific aspect of the dimension, enabling more targeted evaluation. For example, under the ’Logical Flow’ component, attributes such as ”Evaluate whether the connections between different situations are well represented,” ”Check if the descriptions of the same subject are well-linked,” and ”Check if the sentences are related to each other” can be clustered.
@@ 3.4.3. Question generation
In this stage, the clustered attributes from the prior step are transformed into actionable ”Yes-No” questions aimed at probing the identified attributes. First, the LLM generates a key question for each component based on the corresponding attributes. For example, if there are five components, the LLM generates five key questions grounded on each component. Subsequently, the agent breaks down the key questions into detailed sub-questions to more precisely identify the relevant features within a summary. These questions form the core of the checklist, guiding the LLM evaluator in assessing the summaries.
3.4.4. Question validation
In the final stage, generated questions undergo rigorous review by an LLM for clarity, relevance, and effectiveness. The LLM is prompted to combine any redundant questions and drop irrelevant questions, ensuring the checklist is both reliable and practical for final use in rating texts. The questions should be answerable with ”Yes” or ”No” containing concepts of a rating dimension as well. After the validation process, the final questions compose the checklist which is used for text evaluation.
@@ 3.5. Checklist-based Evaluation
As mentioned before, we adopt checklist-based evaluating approach using the checklist created following aforementioned procedures. Therefore, an LLM evaluator is used to directly respond to each checklist question in a binary manner (Yes or No) during this stage. The number of positive responses (‘Yes’ answers) is then tallied. The proportion of positive responses to the total number of questions is scaled from 1.0 to 5.0 as the final score, as shown in Equation 1, to align with the score range of the ground truth scores in the dataset. This method assumes that the ratio of positive responses reflects overall text quality, with each question targeting a specific evaluation component. The scaled scores of each summary are compared with the ground truth scores assigned by humans.
(1) Final Score=Number of Positive ResponsesTotal Number of $\text{Final Score}=\frac{\text{Number of Positive Responses}}{\text{Total Number of Questions}}\times 5.0$
@@ 4. Experiment
The experiment consists of three phases aligned with the research questions. First, we compare our method, InteractEval, with baseline text evaluation methods. Next, we assess the effectiveness of the TA approach in creating checklists for text evaluation. Finally, we explore the impact of combining TA outcomes from humans and LLMs, comparing it to other TA conditions: Single-LLM-TA, Single-Human-TA, Multiple-LLMs-TA, and Multiple-Humans-TA. We employ both quantitative metrics, such as correlation with human evaluations, and qualitative analysis, focusing on the generated attributes.
@@ 4.1. Dataset
We evaluate our InteractEval framework using SummEval (Fabbri et al., 2021), a widely recognized benchmark in the text generation task particularly oriented to news article summarization. SummEval provides source texts, reference texts, and model-generated summaries, along with evaluation dimensions and human scores: Coherence, Fluency, Consistency, and Relevance. To assess our framework effectively and consider its generalizability, we sample 10% of the entire dataset, ensuring that the distribution of human annotations across each dimension and source text is uniformly represented in the sample. This is the way utilized in (Lee et al., 2024b) as well.
@@ 4.2. TA Condition Description
We examine the effectiveness of the TA process in generating attributes for the checklist used in text evaluation. We compare rating performances and qualitatively analyze the text attributes across five different TA conditions: Single-LLM-TA (a single LLM verbalizing its thought process, SL-TA), Single-Human-TA (involving one human expert, SH-TA), Multiple-LLMs-TA (with several LLMs, ML-TA), Multiple-Humans-TA (involving multiple human experts, MH-TA), and Combination-TA (combining human experts and LLMs, Comb-TA).
@@ 4.3.1. Measuring correlation
We measure the performance of InteractEval using sample-level correlation. This method computes correlation for each individual sample based on summaries from multiple source texts and then averages these correlations across all samples (Lee et al., 2024b). For each source text $x_{i}$, where $i\in\{1,2,\dotsc,I\}$, there are $M$ system-generated summaries $y_{i,m}$, with $m\in\{1,2,\dotsc,M\}$. Let $f_{\text{auto}}$ denotes the automatic evaluation metric (e.g., BLEU) and $f_{\text{human}}$ represents the human evaluation scores. The sample-level correlation $C^{\text{sample}}$ for each dimension is defined as:
@@
where $g$ is a correlation coefficient such as Spearman or Kendall’s tau.
@@ 4.3.2. Measuring difference
In this study, the Fisher’s transformation (Fisher, 1970) is employed as a measurement method to normalize and compare the correlation coefficients of the five separate TA conditions. Converting the correlation values into z-scores allows more accurate comparison and analysis of correlations, particularly in hypothesis testing and constructing confidence intervals.
@@ 4.3.3. Measuring similarity
We conduct text analysis to examine the similarity of generated attributes within the same dimension. This analysis evaluates whether each TA condition produces diverse attributes from both a lexical and semantic perspective. For lexical similarity, we use ROUGE (Lin, 2004) and Jaccard similarity (Jaccard, 1912; Niwattanakul et al., 2013), which focus on word overlap and sequences without considering meaning or context. For semantic similarity, we use cosine similarity, which measures semantic closeness rather than exact word matching.

### v2

@@ 4 University Lecturer Teaching English writing in a university Fluent (native)
  • Internal Quality
    • Coherence: The collective quality of all sentences. We align this dimension with the document understanding conference (DUC) (Dang, 2005) quality question of structure and coherence whereby ”the summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to a coherent body of information about a topic.”
    • Fluency: The quality of individual sentences. Drawing again from the DUC (Dang, 2005) quality guidlines, sentences in the summary ”should have no formatting problems, capitalization errors or obviously ungrammatical sentences (e.g., fragments, missing components) that make the text difficult to read.”
  • External Alignment
3.3. Attribute Collection through Think Aloud
In this stage, we collected the text attributes to construct checklists of each dimension, utilizing the TA process, which enhances divergent thinking (Ericsson, 6; Lewis, 1982). The text attributes indicate the key considerations for text evaluation, which are used to design assessment checklists. During the TA process, four human participants verbalized their thoughts while conducting experimental tasks (Ericsson and Simon, 1980). Likewise, four different LLMs articulated the text attributes needed for rating the text quality of each dimension, as shown in Figure 2 (A). To minimize idiosyncratic biases arising from individual members of each participant group, whether humans or LLMs, we included four members in each group. As illustrated in Figure 3, humans engaged in the TA process by reading the definitions of each dimension alongside four sample summaries per dimension, and the LLMs generated text attributes using the same materials provided to the humans. As a result, the human participants and LLMs generated fine-grained attributes by analyzing the detailed features of each dimension as manifested in the example summaries. Between the methods of CTA and RTA (Ericsson, 6), which are two main methods of TA, the current framework used CTA, as it is straightforward to implement (Fan et al., 2019) and free from post-task retrospection biases (Van den Haak et al., 2004; McDonald et al., 2012), with minimal impact on participants’ behavior (Hertzum et al., 2009).
3.3.1. TA by human experts
The human experts consisted of two journalism and two linguistics specialists, chosen based on the nature of the news article summary dataset. We assumed that journalism experts would focus on aspects related to news content while linguistics experts would concentrate on the language features of the texts. Their details are provided in Table 1. The experts conducted the process in real-time via online video calls, addressing each dimension—Coherence, Fluency, Consistency, and Relevance—separately. Specifically, we collected participants’ thoughts on each dimension through the following process. First, participants were assigned an assessor persona and provided with four news articles and summaries, actual scores corresponding to different score ranges, and the rubric of each dimension. Then, participants reviewed the news summaries corresponding to each score range and verbalized their reasoning for why they believe the news articles received the given scores. Finally, the participants’ verbalizations were recorded and transcribed, and the identified text attributes were compiled.
Figure 4. The Overall Process of Text Evaluation: Stage 1: Four humans and four LLMs participate in think-aloud-based text attribute generation, respectively. Stage 2: GPT-4 is solely utilized to construct the checklists. Stage 3: Either GPT-4 or GPT-3.5-Turbo is used to evaluate the quality of the summaries, which are presented with the checklists to guide the evaluation process.
Table 2. System message and predefined template for LLMs’ TA
System Message
You are an assessor who rates {Dimension} of the summary of a news article.
You must read the summary samples and rubric for rating {Dimension}.
While reading, write down your thoughts on what to keep in mind when evaluating the {Dimension} of the text, referring to the rubric.
Avoid mentioning contents specific to the samples.
Instead, focus on general considerations for evaluating the {Dimension} of any summary based on the provided rubric.
Predefined Template
Rubric of {Dimension}:
{Rubric}
<Sample 1>
News Article: {Source Text}
Summary: {Summary}
Score: {Score}
<End of Sample 1>
<Sample 4>
News Article: {Source Text}
Summary: {Summary}
Score: {Score}
<End of Sample 4>
Please provide each consideration in a JSON format with the keys of consideration index and the values of consideration.
Do not name each consideration.
For example, {1: ”consideration 1”, 2: ”consideration 2”, 3: ”consideration 3”, 4: ”consideration 4”, 5: ”consideration 5”, …}
@@ 3.3.2. TA by LLMs
The LLMs included four different types of representative agent: GPT-4 (Achiam et al., 2023), Llama-3.1-8B-Instruct  [^6] https://llama.meta.com/, Claude-3.5-Sonnet  [^7] https://www.anthropic.com/news/claude-3-5-sonnet, and Gemini-1.5-Pro (Reid et al., 2024). The LLMs simulated a human-like TA process using prompt templates based on the TA method as shown in Table 2. In reality, unlike humans, LLMs generate text attributes not through TA but through a process of synthesizing responses to prompts based on their pre-trained background knowledge and text-generation algorithms. However, for ease of exposition, this process is referred to as TA in our paper. The same prompt templates were used across the four mentioned LLMs for attribute generation. For this TA activity, we constructed a prompt template per dimension using the identical materials given to humans during the TA procedure: four news articles and summaries, corresponding scores, and dimension definitions. During the TA process, we assigned the role of an assessor to the LLMs using prompt templates, mirroring the human TA process, where we specifically assigned human participants the assessor role.
@@ 3.4. Checklist Construction
In this stage, we merged the attributes collected from the previous TA process (Section 3.3) of both humans and LLMs. Then, we built the checklists using the combined text attributes. The checklist construction process consists of three steps: 1) Component Extraction, 2) Attribute Clustering, and 3) Question Generation. Each step can utilize any off-the-shelf LLM, such as GPT series, and in our study, we used GPT-4 identically as shown in Figure 4. The integration of attributes from both sources (humans and LLMs) was expected to create a synergistic effect, where the strengths of human expertise and LLMs complement each other, leading to a more refined and effective checklist.
Table 3. System message and predefined template for component extraction
System Message
1. The components should be related to {Dimension} defined below.
2. Remove any components that are ambiguous or redundant.
3. The number of the components should not be more than 5.
Rubric of {Dimension}: {Rubric}
Predefined Template
Attributes: {Attributes}
For example, [”component 1”, ”component 2”, ”component 3”, …]
@@ 3.4.1. Component extraction
The component extraction process identified key themes related to a rating dimension from the text attributes generated in the prior TA process. The components represent key sub-concepts within a dimension, specifying the primary areas of assessment. For instance, a component related to Coherence may involve attributes such as ”Structure”, ”Logical Flow” and ”Tone style.” This step narrows down the focus areas, laying the groundwork for generating key questions tied to these components. Using an LLM as an extractor, principle components were extracted from the merged attributes by prioritizing the most pertinent attributes from both human experts and the LLMs. To avoid noise and redundancy, we prompted the extractor LLM to eliminate ambiguous or repetitive components from the combined attributes. The number of components was limited to five or fewer to prevent excessive subdivision, which could lead to overlapping meanings and create noise. These components then guided the subsequent steps in checklist development. Table 3 represents the prompt template leveraged for extracting components.
@@ 3.4.2. Attributes clustering
After extracting the components from the combined group of attributes, an LLM clustered each attribute under the appropriate component to identify specific associated attributes. This clustering divided the scoring dimension into segmented sections, leveraging the ideas of both humans and LLMs. Each section focused on a specific aspect of the dimension, enabling more targeted evaluation. For example, under the ”Logical Flow” component, attributes such as ”Evaluate whether the connections between different situations are well represented,” ”Check if the descriptions of the same subject are well-linked,” and ”Check if the sentences are related to each other” were clustered. Table 4 shows the prompt template with regard to attributes clustering.
Table 4. System message and predefined template for attributes clustering
System Message
You can remove attributes that are redundant or confusing.
Predefined Template
Components: {Components}
Attributes: {Attributes}
Please, provide your answer in a JSON format with the keys of components and the values of the list of attributes.
For example, {”component 1”: [”attribute 1”, ”attribute 4”], ”component 2”: [”attribute 2”, ”attribute 3”], …}
Table 5. System message and predefined template for question generation
System Message
You are a key question generator for evaluating {Dimension}.
2. Based on these common features, construct a set of Yes-No questions related to each key component.
These questions should be designed to effectively evaluate the accuracy and relevance of a summary of a news article.
Predefined Template
The rubric of {Dimension} and the set of components and their attributes corresponding to {Dimension} are provided below.
Each component and question must satisfy the following conditions:
1. Each component should have one corresponding question.
2. Each question must be answerable with ”Yes” or ”No”.
3. Each question must incorporate concepts from the key component.
5. Formulate questions so that a ”Yes” answer is a positive response.
# Rubric
{Rubric of a dimension}
# Components and attributes
{Components and attributes}
@@ 3.4.3. Question generation
In this stage, the clustered attributes from the prior step were transformed into actionable ”Yes-No” questions aimed at probing the identified attributes. First, the LLM generated a key question for each component based on the corresponding attributes. For example, if there were five components, the LLM generated five key questions grounded on each component. Subsequently, the agent broke down the key questions into detailed sub-questions to more precisely identify the relevant features within a summary. These questions form the core of the checklist, guiding an LLM evaluator in assessing the summaries.
The generated questions underwent a review by a validator LLM to combine any redundant questions and drop irrelevant questions, ensuring a checklist is both reliable and practical for final use in rating texts. The questions should be answerable with ”Yes” or ”No” containing concepts of a rating dimension as well. After the validation process, the final questions composed the checklist that was used for text evaluation. The prompt templates utilized in this stage are shown in Tables 5, 6, and 7. They were used at different phases of the stage. Table 5 details the template for generating questions, Table 6 outlines the template for breaking down these questions into sub-questions, and Table 7 provides the template used for question validation. They were inspired by the prompts introduced in (Lee et al., 2024b). The generated checklist for ”Coherence” is shown in Table 8 (see Appendix A for the checklist of other dimensions).
Table 6. System message and predefined template for sub-question generation
System Message
Predefined Template
In this task, you need to create a question to evaluate the {Dimension} of the summary of the original document.
The rubric of {Dimension} and the questions corresponding to the key component of {Dimension} are provided below.
Use them to generate sub-questions for each key question.
Each sub-question must satisfy the following conditions:
1. Each question must be answerable with ”Yes” or ”No”.
5. Formulate sub-questions so that a ”Yes” answer is a positive response.
6. Each sub-question must focus on asking about the presence of positive aspects rather than the absence of negative ones.
For example, ask whether a summary avoids negative elements rather than contains them.
# Rubric
{Rubric of a dimension}
# Components and corresponding questions
{Components and questions}
Please provide your answer in a JSON format with the keys of components and the values of the list of sub-questions.
For example, {”component 1”: [”sub question 1”, ”sub question 2”,…] , ”component 2”: [”sub question 1”, ”sub question 2”, ”sub question 3”, …],…}
Table 7. System message and predefined template for question validation
System Message
You are a sub-question evaluator.
Predefined Template
In this task, you need to examine each sub-question for evaluating the {Dimension} of the summary of the original document.
The rubric of {Dimension} is provided below.
# Rubric
{Rubric of a dimension}
Each sub-question must satisfy the following conditions:
1. Each question must be answerable with ”Yes” or ”No”.
2. Each question must contain concepts of the {Dimension}.
5. Formulate questions so that a ”Yes” answer is a positive answer.
7. Each sub-question must focus on asking about the presence of positive aspects rather than the absence of negative ones.
For example, ask whether a summary avoids negative elements rather than contains them.
# Sub-questions
{Sub-questions}
For example, [”sub question 1”, ”sub question 2”, ”sub question 3”, …]
Table 8. Checklist of Coherence dimension generated via the checklist construction process
Dimension Questions
Coherence - Does the summary present information in a logical sequence?
- Are there smooth transitions between sentences and topics in the summary?
- Does the summary avoid redundancy or repetition that could disrupt the flow of information?
- Does each sentence in the summary contribute to the overall structure or flow of ideas?
- Are pronouns and references used in the summary clear and consistent?
- Are key actors, events, and relationships introduced clearly in the summary?
- Is the level of detail maintained consistently throughout the summary?
- Does the summary avoid sudden shifts in the level of detail?
- Does the summary provide a balanced amount of detail for all key points?
@@ 3.5. Checklist-based Evaluation
As mentioned before, we adopt checklist-based evaluation approach using the checklist created following aforementioned procedures. Therefore, an LLM evaluator is used to directly respond to each checklist question in a binary manner (Yes or No) during this stage. Any LLM can be used at this stage. In our case, as shown in Figure 4, we use GPT-3.5-Turbo as the evaluator when comparing our performance with the baselines that use GPT-3.5-Turbo, and we use GPT-4 as the evaluator when comparing with baselines that leverage GPT-4. The number of positive responses (‘Yes’ answers) is then tallied. The proportion of positive responses from the evaluator to the total number of questions is scaled from 1.0 to 5.0 as the final score, to be aligned with the score range of the ground truth scores in the dataset. This method assumes that the ratio of positive responses reflects overall text quality, with each question targeting a specific evaluation component. The scaled scores of each summary are compared with the ground truth scores assigned by humans, which are provided in the dataset.
@@ 4. Experiment
Figure 5. Comparison of Distributions across Four Dimensions in SummEval Dataset: Entire Dataset refers to the distribution of the entire dataset with respect to each dimension, while 1st Sampled Dataset and 2nd Sampled Dataset represent the distributions of subsets obtained by sampling 10% of the entire dataset.
The experiment consists of three phases aligned with the research questions. First, we compare our method, InteractEval, with baseline text evaluation methods. Next, we assess the effectiveness of the TA approach in creating checklists for text evaluation. Finally, we explore the impact of combining TA outcomes from humans and LLMs, comparing it with other TA conditions: Single-LLM-TA, Single-Human-TA, Multiple-LLMs-TA, and Multiple-Humans-TA. We employ both quantitative metrics, such as correlation with ground-truth scores, and qualitative analysis, focusing on the generated attributes. The code is available at Github repository  [^8] https://github.com/BBeeChu/InteractEval.git.
@@ 4.1. Dataset
We mainly evaluated our InteractEval framework on SummEval dataset (Fabbri et al., 2021), a widely recognized benchmark in the text generation task particularly oriented to news article summarization. SummEval contains 1,600 data instances providing source texts, reference texts, and model-generated summaries, evaluated on four dimensions: Coherence, Fluency, Consistency, and Relevance. Each dimension was independently scored by three human annotators using a scale ranging from 1.0 to 5.0, with the final score for each dimension determined by averaging their individual ratings. Additionally, to examine the generalizability of InteractEval’s performance on a different text evaluation task, we utilized the ELLIPSE dataset (Crossley et al., 2023), comprising 6,482 English essays written by non-native ESL (English as a second language) learners. Each essay was scored on seven dimensions: Overall, Cohesion, Conventions, Grammar, Phraseology, Syntax, and Vocabulary. These dimensions were respectively evaluated by two human assessors on a scale of 1.0 to 5.0, and the final score was calculated as the average of their ratings.
@@ 4.2. TA Condition Description
We examine the effectiveness of the TA process in generating attributes for the checklist used in text evaluation. We compare rating performances and qualitatively analyze the text attributes across five different TA conditions: Single-LLM-TA (involving a single LLM articulating text attributes, hereafter referred to as SL-TA), Single-Human-TA (involving one human expert verbalizing one’s thought process, SH-TA), Multiple-LLMs-TA (involving several LLMs articulating text attributes independently, ML-TA), Multiple-Humans-TA (involving multiple human experts verbalizing their thought process independently, MH-TA), and Combination-TA (combining human experts’ and LLMs’ TA outcomes, Comb-TA).
@@ 4.3.1. Measuring correlation
We measure the performance of InteractEval using sample-level correlation, a method widely adopted in prior LLM-based studies (Lee et al., 2024b; Fu et al., 2023; Chen et al., 2023a; Ahia et al., 2023; Wang et al., 2024a). This approach helps reduce evaluation costs while maintaining effective performance assessment. Although prior studies have noted the high cost problem of using GPT-series models with large datasets (Chen et al., 2023b; Ong et al., 2024), we believe that instance-level evaluation in real-world applications can alleviate this concern. Therefore, we continue using GPT-series models, which are a leading choice for text evaluation. This method computes correlation for each individual sample based on summaries from multiple source texts and then averages these correlations across all samples (Lee et al., 2024b). For each source text $x_{i}$, where $i\in\{1,2,\dotsc,I\}$, there are $M$ system-generated summaries $y_{i,m}$, with $m\in\{1,2,\dotsc,M\}$. Let $f_{\text{auto}}$ denote the automatic evaluation metric (e.g., BLEU) and $f_{\text{human}}$ represent the human evaluation scores. The sample-level correlation $C^{\text{sample}}$ for each dimension is defined as:
@@
where $g$ is a correlation coefficient such as Spearman’s Rho ($\rho$) or Kendall’s Tau ($\tau$). To ensure the representativeness of sample-level correlations, we sampled 10% data instances from the original dataset while preserving the identical distribution of human annotations (ground truth scores) across each dimension as found in the original one. Figure 5 demonstrates that the score distributions of the entire SummEval dataset and the sampled subsets for each dimension are nearly identical, confirming their representativeness (see Appendix D for the score distributions in ELLIPSE dataset). Additionally, we evaluated our method’s performance against baselines in two trials to assure the robustness of our framework.
@@ 4.3.2. Measuring difference
In this study, the Fisher’s transformation (Fisher, 1970) is employed to normalize and compare the correlation coefficients of the five separate TA conditions. Converting the correlation values into z-scores allows more accurate comparison and analysis of correlations, particularly in hypothesis testing and constructing confidence intervals.
@@ 4.3.3. Measuring similarity
We conduct text analysis to examine the similarity of generated attributes within the same dimension. This analysis evaluates whether each TA condition produces diverse attributes from both a lexical and a semantic perspective. For lexical similarity, we use ROUGE (Lin, 2004) and Jaccard similarity (Jaccard, 1912; Niwattanakul et al., 2013), which focus on overlapping words or sequences without considering meaning or context. For semantic similarity, we use cosine similarity, which measures semantic closeness rather than exact word matching.